# The Dawn of Natural Language to SQL: Are We Fully Ready?

### v1

@@ Abstract.
The source code, data, and/or other artifacts have been made available at https://github.com/BugMaker-Boyan/NL2SQL360/.
@@ 1. Introduction
Natural Language to SQL (nl2sql), which converts a natural language query (nl) into an SQL query (sql), can significantly lower the barrier for both lay users and expert users in accessing massive datasets and deriving insights (Gu et al., 2023; Chen et al., 2023; Wang et al., 2022; Liu et al., 2022; Pourreza and Rafiei, 2023; Gao et al., 2023; Li et al., 2023d). Especially being empowered by the recent advances of large language models, the performance of nl2sql solutions has been significantly improved. The trend of providing nl2sql solutions by database vendors has shifted from a myth to must-go.
@@
Q1: Where Are We Now? Figure 1 depicts the evolution of nl2sql methods in the last two decades, from rule-based methods, deep neural network-based methods, tunable pre-trained language models (PLMs), to giant large language models (LLMs), alongside the development of benchmarks like Spider (Yu et al., 2018) and BIRD (Li et al., 2023c). Note that LLMs (e.g., GPT-4 (OpenAI et al., 2024) and Llama2 (Touvron et al., 2023)) are larger language models compared to PLMs (e.g., GPT-2 (Radford et al., 2019) and BART (Lewis et al., 2019)) and exhibit advanced language understanding and emergent abilities (Minaee et al., 2024). Employing PLMs for the nl2sql task requires fine-tuning on task-specific datasets, while harnessing LLMs for this task can be done through prompts (in-context learning) for all kinds of LLMs or fine-tuning (i.e., instruction following) for open-source LLMs only (Zhao et al., 2023). State-of-the-art (SOTA) results are achieved by both PLM- and LLM-based methods.
Figure 1. An Overview of nl2sql Methods
Figure 2. Evolution of PLM- and LLM-based nl2sql Models
[…]
@@
[Various Data Domains.] BI platforms like Tableau (tab, [n.d.]) often have various database domains (e.g., movies and sports) with unique schemas and terminologies. An ideal nl2sql model must generalize across these varied domains while adapting to each specific domain to meet ad-hoc requirements effectively.
@@
[New Linguistic Phenomena.] For the same query intent, different users may pose nl questions with different abbreviations, synonyms, and question styles. Thus, the ability of an nl2sql model to comprehend and accurately interpret a wide spectrum of nl query variants becomes important.
Let’s better illustrate the comparisons of different nl2sql models from different use cases using empirical results.
@@
[[[[Various Data Domains]]]] Figure 3(a) compares different models in the Competition domain. The result shows that fine-tuning-based LLM/PLM methods outperform all prompt-based LLM methods. Specifically, the best PLM-based method, RESDSQL-3B$+$NatSQL (Li et al., 2023d), achieves $83.9\%$ execution accuracy, which outperforms the best prompt-based LLM method, DAILSQL (with GPT-4) (Gao et al., 2023), by $3.3\%$. The above observations suggest that fine-tuning is a crucial strategy for enhancing the domain adaptation capabilities of nl2sql models.
@@
Q3: Can we combine the best of both worlds and design a super NL2SQL model? The question following Q1 and Q2 is: if there is no single winner on different scenarios, can we design a super nl2sql model that combines the merits of both PLMs and LLMs and is robust for different scenarios.
Contributions. In this paper, we systematically evaluate different PLM- and LLM-based nl2sql models on different benchmarks, from different angles. During these extensive experiments, we built a testbed that can help researchers and practitioners better evaluate nl2sql models on their specific scenarios, observed interesting experimental findings, and designed a super nl2sql model that is the most robust than SOTA solutions.
Our main contributions are summarized as follows.
@@
N-best Rerankers + PICARD (Zeng et al., 2023) T5 N/A PICARD N-best Rerankers
@@
SHiP + PICARD (Zhao et al., 2022) T5 N/A PICARD
@@
(2) New experimental findings. We tested 13 LLM-based and 7 PLM-based nl2sql solutions on the Spider and BIRD datasets, varying 15 different settings to analyze their performance in various usage scenarios (Section 4). The key findings are as follows:
(i) Accuracy. Fine-tuning is crucial for enhancing performance. Specifically, LLM-based methods with fine-tuning excel in the EX metric, while PLM-based methods lead in the EM metric. However, they can be distinguished as winners in subsets of sql with specific characteristics. For example, methods using GPT-4 perform notably better with subqueries.
(iii) Domain Adaption. For nl2sql tasks across different domains, there is no clear winner between LLM-based and PLM-based methods. However, in-domain data during fine-tuning process is crucial for model performance in specific domains.
2. Natural Language to SQL
2.1. Related Works: A Bird’s-Eye View
Rule-based Methods. We can observe that early work was primarily based on pre-defined rules or semantic parsers (Rajkumar et al., 2022; Li and Jagadish, 2014; Katsogiannis-Meimarakis and Koutrika, 2023, 2021). For example, NaLIR (Li and Jagadish, 2014) employs a syntactic parser to understand the nl query and links to the database elements and then relies on handcrafted rules to generate the sql query. However, these methods exhibit significant limitations in terms of their adaptability, scalability, and ability to generalize.
Neural Network-based Methods. To overcome these limitations, researchers have turned to employing neural networks to learn the translation from nl queries to sql queries. During this period, numerous large-scale benchmark datasets were released, including WikiSQL (Zhong et al., 2017), Spider (Yu et al., 2018), etc. In this line of research, sequence-to-sequence based nl2sql methods (Zhong et al., 2017; Cheng et al., 2017; Xiao et al., 2016) were developed and reached a new bar at that time. For example, IRNet (Guo et al., 2019) utilizes an encoder to encode the nl query and database schema and then uses a decoder network to generate the sql query.
PLM-based Methods. Around 2017, with the introduction of the Transformer (Vaswani et al., 2017) and the Spider dataset, methods based on neural networks began to emerge, quickly becoming the mainstream approach. The advent of models like BERT (Devlin et al., 2018) and T5 (Raffel et al., 2020) marked the rise of pre-trained language models-based methods (Li et al., 2023b, d; Scholak et al., 2021), which achieved competitive results on benchmark datasets. For example, RESDSQL (Li et al., 2023d) utilizes a two-stage framework for nl2sql. First, it identifies relevant schema elements such as table names and columns directly from the natural language query. Then, it uses these elements to construct the SQL query, which is one of the best-ranked models in the Spider leaderboard.
LLM-based Methods. Recently, the emergence of giant large language models like ChatGPT and GPT-4 (Achiam et al., 2023) has led to a new wave of solutions. These LLM-based nl2sql methods have become the most prominent and representative solutions in the current nl2sql landscape (Pourreza and Rafiei, 2023; Gao et al., 2023; Dong et al., 2023; Li et al., 2024; Wang et al., 2023). For example, DAIL-SQL (Gao et al., 2023) leverages GPT-4 through effective prompt engineering methods, achieving competitive results on the Spider dataset.
We can observe that the schema linking module, a component integral to most approaches, highlights its crucial role in the nl2sql process. Furthermore, the incorporation of database content into all PLM-based methods signifies its essential contribution to enhancing the accuracy and relevance of the generated sql queries. This emphasizes the foundational importance of understanding both the schema and content of databases for the nl2sql task. In the sql generation step, all PLM-based methods adopt beam search-like strategies, e.g., PICARD (Scholak et al., 2021), to identify the optimal output tokens within the constraints imposed by sql syntax rules. Conversely, LLM-based methods rely on greedy-based strategies for sql generation. In the post-processing step, most LLM-based methods incorporate heuristic prompting strategies, such as Self-Correction and Self-Consistency, to refine the initial outputs, ensuring they align more closely with the intended sql queries.
@@ 2.2. Existing Experiments and Their Limitations
Existing Experiments. There are several experimental studies relevant to our research. For example, Gao et al. (Gao et al., 2023) evaluated the potential of open-source LLMs for nl2sql tasks through prompt engineering. Rajkumar et al. (Rajkumar et al., 2022) explored the capabilities of the Codex language model in handling the nl2sql task under zero-shot and few-shot settings. Gkini et al. (Gkini et al., 2021) conducted an in-depth evaluation of parsing-based and keyword-based nl2sql. While the first two studies mainly focused on evaluating LLM-based nl2sql solutions, the third investigated parsing-based nl2sql methods.
@@
(3) Limited Exploration of the NL2SQL Design Space. A gap in the current nl2sql research and practice is the limited exploration of the design space of the nl2sql framework based on both LLM and PLM approaches. This lack of comprehensive research restricts our understanding of how different architectural and functional modules from both LLM and PLM can be synergistically incorporated to enhance nl2sql systems.
@@ 3. NL2SQL360: A Testbed for NL2SQL
Figure 4. An Overview of NL2SQL360
We design a testbed, as shown in Figure 4, for evaluating and analyzing nl2sql solutions. NL2SQL360 can help researchers learn the design choices in nl2sql systems and compare SOTA models with less development effort, and provides practitioners with experimental findings of different types of models in specific scenarios to promote practical applications.
Figure 4 overviews our testbed framework, comprising six core components: datasets repository, model zoo, dataset filter, evaluated metrics, nl2sql evaluator, and analysis module.
@@
Dataset Filter. Traditional evaluations, averaging performance across entire benchmark datasets, miss nuanced nl2sql performance insights for varied scenarios outlined in Section 1. To tailor evaluations to distinct scenarios, we select specific subsets of benchmarks, including particular databases, nl, and sql queries. These subsets highlight unique traits, such as query complexity, database schema diversity, and distinctive sql features like JOIN operations or nested queries. Therefore, we introduce a dataset filtering mechanism in our NL2SQL360. This allows for the segregation of testing datasets into more focused subsets based on various criteria:
@@
(1) $\small QVT=\frac{1}{M}\sum_{i=1}^{M}\left(\frac{\sum_{j=1}^{m_{i}}\mathbbm{1}\left(\mathcal{F}(N_{ij})=Q_{i}\right)}{m_{i}}\right)$
@@
  • $\mathbbm{1}(\cdot)$ is the indicator function that returns 1 if the query results inside are equal, and 0 otherwise.
@@ 4.1. Experimental Settings
(2) DAILSQL (Gao et al., 2023) encodes the question and database schema in sql code style. It selects few-shot examples based on their structural (skeleton) similarities and query similarities. These elements are combined into an efficient prompt to guide GPT-4.
(3) DAILSQL(SC) (Gao et al., 2023) is the version of DAILSQL with a Self-Consistency (SC) strategy for post-processing.
@@
(10) Llama3-8B (AI@Meta, 2024) on over 15T token of data – a training dataset 7x larger than that used for Llama 2, including 4x more code.
@@ 4.2. Experiments on Evaluating Accuracy
Exp-1: Overall Accuracy on Benchmarks. We evaluate the performance of LLM-based and PLM-based methods across sql queries of different complexities. We run all methods on the Spider and BIRD development sets and compute their Execution Accuracy (EX) and Exact Match Accuracy (EM) metrics. Note that we retrained the official PLM-based method RESDSQL from scratch on the BIRD train set. Since the complete code for NatSQL was not publicly available, our models did not incorporate NatSQL. Additionally, due to constraints on GPT’s resources, we did not reproduce the DINSQL method on BIRD. Table 3 and Table 4 report the results.
The state-of-the-art (SOTA) EX and EM in specific SQL complexity are marked as orange and blue in the table, respectively.
@@
Figure 6. EX vs. SQL Characteristics on Spider.
Figure 7. EX vs. SQL Characteristics on BIRD.
@@
To this end, we classify sql queries based on four criteria: (1) the presence of subqueries, (2) the number of logical connectors, (3) the use of ORDER BY, and (4) and the number of JOINs. Note that our NL2SQL360 supports sql query filtering based on individual sql clauses, their combinations, or user-defined conditions. However, due to space constraints, we demonstrate only four representative aspects. We run all methods on these four subsets of sql queries and compute their EX metrics.
We further classify LLM-based methods into prompt-based and fine-tuning-based LLMs. Figure 5 visualizes the EX performance distribution across different subsets of the Spider and BIRD datasets. Figure 7 and Figure 7 further show detailed results of various methods across different subsets. The bar chart shows the overall EX for each method. In the heatmap, the x-axis represents different methods, and the y-axis represents different subsets.
@@
In scenarios without Logical Connectors, as shown in Figure 5, LLM-based methods show no significant advantage over PLM-based methods on the Spider dataset. However, on the BIRD dataset, LLM-based methods outperform PLM-based methods, possibly due to the higher complexity of the BIRD dataset, as indicated in Table 2. In scenarios requiring Logical Connectors, the LLM-based methods consistently outperform PLM-based methods on both datasets.
@@
SQL with JOIN. However, for scenarios requiring JOIN operations, LLM-based methods outperform PLM-based methods on both datasets. This could be due to the JOIN operation’s need for understanding complex database schemas, where LLMs typically excel due to their superior context understanding capabilities.
@@
Exp-2.4: #-ORDER BY. As shown in Figure 5, we observed that without ORDER BY clause, LLM-based methods outperform PLM-based methods on both the Spider and BIRD datasets. However, with the ORDER BY clause, LLM-based methods underperform compared to PLM-based methods on the Spider dataset, while they outperform PLM-based methods on the BIRD dataset. This difference might be because the BIRD dataset is more complex than the Spider dataset.
Finding 5. In scenarios that include the ORDER BY clause, the performance of LLM-based and PLM-based methods varies across different datasets. In general, LLM-based methods demonstrate stronger generalization capability.
@@
Finding 6. There is no clear winner between LLM-based methods and PLM-based methods in QVT. Fine-tuning the model with task-specific datasets may help stabilize its performance against nl variations.
@@
Finding 7. Different methods exhibit varying biases towards different domains, and there is no clear winner between LLM-based and PLM-based methods. However, in-domain training data during fine-tuning process is crucial for model performance in specific domains.
@@
Exp-5: Supervised Fine-tuning on LLM-based Methods. We investigated Supervised Fine-tuning (SFT) of open-source LLMs for the nl2sql task. DAILSQL (Gao et al., 2023) examines the impact of varying shot and prompt representation during SFT but does not address which open-source LLMs are best suited for SFT in the nl2sql task. DAILSQL found that SQL-style prompts were beneficial, so we adopted a similar prompt approach in a zero-shot setting, as shown in Figure 10. Given that nl2sql is a code-related task, we selected five open-source LLMs with varying code abilities, evaluated using the HumanEval (Pass@1) metric (Chen et al., 2021). To ensure a fair comparison and account for hardware limitations, all chosen LLMs have similar parameters. The suffix in the model name, such as 7B, indicates the model has 7 billion parameters.

### v2

@@ Abstract.
PVLDB Reference Format:
Boyan Li, Yuyu Luo, Chengliang Chai, Guoliang Li, Nan Tang. PVLDB, 17(11): XXX-XXX, 2024.
doi:XX.XX/XXX.XX Yuyu Luo is the corresponding author.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 11 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/HKUSTDial/NL2SQL360.
@@ 1. Introduction
Figure 1. An Overview of nl2sql Methods.
Figure 2. Evolution of PLM- and LLM-based nl2sql Models.
[…]
Natural Language to SQL (nl2sql), which converts a natural language query (nl) into an SQL query (sql), can significantly lower the barrier for both lay users and expert users in accessing massive datasets and deriving insights (Gu et al., 2023; Chen et al., 2023; Wang et al., 2022; Liu et al., 2022; Pourreza and Rafiei, 2023; Gao et al., 2024; Li et al., 2023d; Luo et al., 2018b, a, 2021, 2022b; Tang et al., 2022; Xie et al., 2024; Luo et al., 2022a; Shen et al., 2021; Liu et al., 2024). Especially being empowered by the recent advances of large language models, the performance of nl2sql solutions has been significantly improved. The trend of providing nl2sql solutions by database vendors has shifted from a myth to a must-go.
@@
Q1: Where Are We Now? Figure 1 shows the evolution of nl2sql methods in the last two decades, from rule-based methods, deep neural network-based methods, tunable pre-trained language models (PLMs), to giant large language models (LLMs), alongside the development of benchmarks like Spider (Yu et al., 2018) and BIRD (Li et al., 2023c). Note that LLMs (e.g., GPT-4 (OpenAI, 2023) and Llama2 (Touvron et al., 2023)) are larger language models compared to PLMs (e.g., GPT-2 (Radford et al., 2019) and BART (Lewis et al., 2020)) and exhibit advanced language understanding and emergent abilities (Minaee et al., 2024; Tang et al., 2024; Ye et al., 2024; Zhu et al., 2024). Employing PLMs for the nl2sql task requires fine-tuning on task-specific datasets, while harnessing LLMs for this task can be done through prompts (in-context learning) for all kinds of LLMs or fine-tuning (i.e., instruction following) for open-source LLMs only (Zhao et al., 2023). State-of-the-art (SOTA) results are achieved by both PLM- and LLM-based methods.
@@
[Various Data Domains.] BI platforms like Tableau (tab, 2024) often have various database domains (e.g., movies and sports) with unique schemas and terminologies. An ideal nl2sql model must generalize across these varied domains while adapting to each specific domain to meet ad-hoc requirements effectively.
@@
[New Linguistic Phenomena.] For the same query intent, different users may pose nl questions with different abbreviations, synonyms, and question styles. Thus, the ability of an nl2sql model to accurately interpret various nl query variants is crucial.
@@
[[[[Various Data Domains]]]] Figure 3(a) compares different models in the Competition domain. The result shows that fine-tuning-based LLM/PLM methods outperform all prompt-based LLM methods. Specifically, the best PLM-based method, RESDSQL-3B$+$NatSQL (Li et al., 2023d), achieves $83.9\%$ execution accuracy, which outperforms the best prompt-based LLM method, DAILSQL (with GPT-4) (Gao et al., 2024), by $3.3\%$. The above observations suggest that fine-tuning is a crucial strategy for enhancing the domain adaptation capabilities of nl2sql models.
@@
Q3: Can we combine the best of both worlds and design a super NL2SQL model? The question following Q1 and Q2 is: if there is no single winner in different scenarios, can we design a super nl2sql model that combines the merits of both PLMs and LLMs and is robust for different scenarios.
Contributions. In this paper, we present NL2SQL360, a testbed designed to systematically evaluate PLM- and LLM-based nl2sql models across different benchmarks from multiple perspectives. NL2SQL360 can help researchers and practitioners better evaluate nl2sql models on their specific scenarios, uncover insightful experimental findings, and design a superior nl2sql model that is more robust than SOTA solutions. Our contributions are as follows.
@@
(2) New experimental findings. We tested 13 LLM-based and 7 PLM-based nl2sql solutions on the Spider and BIRD datasets, varying in 15 different settings to analyze their performance in various usage scenarios (Section 4). The key findings are as follows:
(i) Accuracy. Fine-tuning is crucial for enhancing performance. Specifically, LLM-based methods with fine-tuning excel in the EX metric, while PLM-based methods lead in the EM metric. Furthermore, PLM/LLM-based methods can be distinguished as winners in subsets of sql with specific characteristics. For example, methods using GPT-4 perform notably better with subqueries.
(iii) Domain Adaption. For nl2sql tasks across different domains, there is no clear winner between LLM-based and PLM-based methods. However, in-domain data during the fine-tuning process is crucial for model performance in specific domains.
2. Natural Language to SQL
2.1. Related Works: A Bird’s-Eye View
Rule-based Methods. Early work relied on pre-defined rules or semantic parsers (Rajkumar et al., 2022; Li and Jagadish, 2014; Katsogiannis-Meimarakis and Koutrika, 2023, 2021). For example, NaLIR (Li and Jagadish, 2014) uses a syntactic parser and handcrafted rules to convert nl queries into sql queries. However, these methods are limited in adaptability, scalability, and generalization.
Neural Network-based Methods. To address these limitations, researchers began using neural networks to translate nl queries to sql queries. Several large-scale benchmark datasets such as WikiSQL (Zhong et al., 2017) and Spider (Yu et al., 2018) were released. Sequence-to-sequence nl2sql methods such as IRNet (Guo et al., 2019) were developed. IRNet encodes nl queries and database schemas with an encoder and generates sql queries with a decoder.
PLM-based Methods. Around 2017, the introduction of the Transformer (Vaswani et al., 2017) and the Spider dataset led to the rise of neural network-based methods, which soon dominated the field. Models like BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020) heralded the era of pre-trained language models, achieving top results on benchmark datasets (Li et al., 2023b, d; Scholak et al., 2021). For instance, RESDSQL (Li et al., 2023d), a top performer on the Spider leaderboard, uses a two-stage framework: it first identifies relevant schema elements from the nl query, then constructs the sql query.
LLM-based Methods. The emergence of large language models like ChatGPT and GPT-4 (OpenAI, 2023) has revolutionized nl2sql solutions. LLM-based methods now dominate the nl2sql landscape (Pourreza and Rafiei, 2023; Gao et al., 2024; Dong et al., 2023; Li et al., 2024; Wang et al., 2023). For instance, DAIL-SQL (Gao et al., 2024) uses GPT-4 with prompt engineering, achieving competitive results on the Spider dataset.
The schema linking module is vital to the nl2sql process, as it enhances the accuracy and relevance of generated sql queries by incorporating database content. This underscores the importance of understanding both the schema and content of databases for nl2sql tasks. During sql generation, PLM-based methods use beam search-like strategies (e.g., PICARD (Scholak et al., 2021)) to find optimal tokens within sql syntax constraints, while LLM-based methods use greedy strategies. For post-processing, LLM-based methods employ heuristic prompting strategies like Self-Correction and Self-Consistency to refine outputs to better match intended sql queries.
@@ 2.1. Related Works: A Bird’s-Eye View
N-best Rerankers + PICARD (Zeng et al., 2022) T5 N/A PICARD N-best Rerankers
@@
SHiP + PICARD (Hu et al., 2023) T5 N/A PICARD
@@ 2.2. Existing Experiments and Their Limitations
Existing Experiments. There are several experimental studies relevant to our research. For example, Gao et al. (Gao et al., 2024) evaluated the potential of open-source LLMs for nl2sql tasks through prompt engineering. Rajkumar et al. (Rajkumar et al., 2022) explored the capabilities of the Codex language model in handling the nl2sql task under zero-shot and few-shot settings. Gkini et al. (Gkini et al., 2021) conducted an in-depth evaluation of parsing-based and keyword-based nl2sql. While the first two studies mainly focused on evaluating LLM-based nl2sql solutions, the third investigated parsing-based nl2sql methods.
@@
(3) Limited Exploration of the NL2SQL Design Space. Current nl2sql research and practice have limited exploration of the nl2sql framework’s design space using both LLM and PLM-based modules. This restricts our understanding of how different architectural and functional modules from both LLM and PLM can be synergistically incorporated to enhance nl2sql solutions.
@@ 3. NL2SQL360: A Testbed for NL2SQL
Figure 4 shows the framework of our testbed, NL2SQL360, comprising six core components, as discussed below.
Figure 4. An Overview of NL2SQL360.
@@
Dataset Filter. Traditional evaluations, which average performance across whole benchmark datasets, miss nuanced nl2sql insights for different scenarios (Section 1). To address this, we select specific benchmark subsets, including particular databases, nl, and sql queries, highlighting unique traits like query complexity, database schema diversity, and sql features such as JOIN operations or nested queries.
Therefore, we introduce a dataset filtering mechanism in our NL2SQL360. This allows for the segregation of testing datasets into more focused subsets based on various criteria:
@@
(1) $\small QVT=\frac{1}{M}\sum_{i=1}^{M}\left(\frac{\sum_{j=1}^{m_{i}}\mathds{1}\left(\mathcal{F}(N_{ij})=Q_{i}\right)}{m_{i}}\right)$
@@
  • $\mathds{1}(\cdot)$ is the indicator function that returns 1 if the query results inside are equal, and 0 otherwise.
@@ 4.1. Experimental Settings
(2) DAILSQL (Gao et al., 2024) encodes the question and database schema in sql code style. It selects few-shot examples based on their structural (skeleton) similarities and query similarities. These elements are combined into an efficient prompt to guide GPT-4.
(3) DAILSQL(SC) (Gao et al., 2024) is the version of DAILSQL with a Self-Consistency (SC) strategy for post-processing.
@@
(10) Llama3-8B (AI@Meta, 2024) on over 15T tokens of data – a training dataset 7x larger than that used for Llama 2, including 4x more code.
@@ 4.2. Experiments on Evaluating Accuracy
Exp-1: Accuracy vs. SQL Complexity. We evaluated all methods by varying sql complexity using Spider and BIRD development sets. We computed Execution Accuracy (EX) and Exact Match Accuracy (EM) metrics. Note that we retrained RESDSQL from scratch on the BIRD train set but did not include NatSQL due to unavailable code. Additionally, We omitted DINSQL due to GPT resource constraints.
Table 3 and Table 4 report the results. The state-of-the-art (SOTA) EX and EM in specific SQL complexity are marked as orange and blue in the table, respectively.
@@
Figure 6. EX vs. SQL Characteristics on Spider.
Figure 7. EX vs. SQL Characteristics on BIRD.
@@
To this end, we classify sql queries based on four criteria: (1) the presence of subqueries, (2) the number of logical connectors, (3) the use of ORDER BY, and (4) the number of JOINs. Note that our NL2SQL360 supports sql query filtering based on individual sql clauses, their combinations, or user-defined conditions. However, due to space constraints, we demonstrate only four representative aspects. We run all methods on these four subsets of sql queries and compute their EX metrics.
We classify LLM-based methods into prompt-based and fine-tuning-based. Figure 5 shows the EX performance distribution across subsets of the Spider and BIRD datasets. Figure 7 and  7 provide detailed results. The bar chart presents the overall EX for each method, with methods on the x-axis and subsets on the y-axis.
@@
Without Logical Connectors (Figure 5), LLM-based methods do not outperform PLM-based methods on the Spider dataset. However, on the more complex BIRD dataset (Table 2), LLM-based methods excel. When Logical Connectors are needed, LLM-based methods consistently outperform PLM-based methods on both datasets.
@@
SQL with JOIN. However, for scenarios requiring JOIN operations, LLM-based methods outperform PLM-based methods in both datasets. This could be due to the JOIN operation’s need for understanding complex database schemas, where LLMs typically excel due to their superior context-understanding capabilities.
@@
Exp-2.4: #-ORDER BY. As shown in Figure 5, we observed that without the ORDER BY clause, LLM-based methods outperform PLM-based methods on both the Spider and BIRD datasets. However, with the ORDER BY clause, LLM-based methods underperform compared to PLM-based methods on the Spider dataset, while they outperform PLM-based methods on the BIRD dataset. This difference might be because the BIRD dataset is more complex than the Spider dataset.
Finding 5. In scenarios with the ORDER BY clause, the performance of PLM/LLM-based methods varies across different datasets. Generally, LLM-based methods demonstrate stronger generalization capability.
@@
Finding 6. There is no clear winner between LLM-based and PLM-based methods in QVT. Fine-tuning the model with task-specific datasets may help stabilize its performance against nl variations.
@@
Finding 7. Different methods exhibit varying biases towards different domains, and there is no clear winner between LLM-based and PLM-based methods. However, in-domain training data during fine-tuning is crucial for model performance in specific domains.
@@
Exp-5: Supervised Fine-tuning of LLM-based Methods. We investigated Supervised Fine-tuning (SFT) of open-source LLMs for the nl2sql task. DAILSQL (Gao et al., 2024) examines the impact of varying shot and prompt representation during SFT but does not address which open-source LLMs are best suited for SFT in the nl2sql task. DAILSQL found that SQL-style prompts were beneficial, so we adopted a similar prompt approach in a zero-shot setting, as shown in Figure 10. Given that nl2sql is a code-related task, we selected five open-source LLMs with varying code abilities, evaluated using the HumanEval (Pass@1) metric (Chen et al., 2021). To ensure a fair comparison and account for hardware limitations, all chosen LLMs have similar parameters. The suffix in the model name, such as 7B, indicates the model has 7 billion parameters.